OpenStudy (anonymous):

A 40-lb container of peat moss measures 14x20x30 in. A 40-lb container of topsoil has a volume of 1.9 gal. (a) Calculate the average densities of peat mass and topsoil in units of g/cm3. Would it be correct to say that peat moss is "lighter" than topsoil? Explain. (b) How many bags of the peat moss are needed to cover an area measuring 10 ft. by 20 ft. to a depth of 20 in.?

OpenStudy (anonymous):

@BarbaraKara Please help me with this problem. :3

OpenStudy (anonymous):

My letter (A) is 0.1 g/cm3. I'm not really sure about my answer.

OpenStudy (anonymous):

is lb pounds?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay, to be honest, i have no idea how to convert these measurements - im only used to metric! im really sorry!

OpenStudy (anonymous):

Ohh. Okay, it's alright. Thanks for the help ^_^

OpenStudy (anonymous):

@Festinger : please help me with this. :3

OpenStudy (festinger):

i don't like to work with pounds and inches too. oh and feet. sorry.

OpenStudy (festinger):

Just cancel off the units by multiplying 1: as an example I want to convert km/h to m/s 1m=1000km 1h=3600s so: \[\frac{km}{h}=\frac{\cancel{km}}{h}*1*1=\frac{km}{\cancel{h}}*\frac{1000m}{\cancel{km}}*\frac{\cancel{1h}}{3600s}=\] \[Density=\frac{mass}{volume}\] use google. they provide unit convertors fr you to check your answers!